OpenStudy (bbb911):

Solve by the elimination method. x + y = -2, 4x - 6y = -3

OpenStudy (nincompoop):

multiply the first equation by -4 so you can eliminate x

OpenStudy (anonymous):

Okay in order to use this method we need to multiply the top equation by -4 -4(x+y=-2) -4x-4y=8 4x-6y=-3 now eliminate x -10y=5

OpenStudy (anonymous):

now you can solve for the y coordinate, and plug that back in to find the x coordinate

OpenStudy (nincompoop):

y variable :)

OpenStudy (anonymous):

y-variable * thank you @nincompoop

OpenStudy (bbb911):

sorry i took long thank you guys :)